• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Yi Qiang

    Log entry time 00:27:49 on April23,2008

    Entry number 233838

    keyword=Swing Shift Summary

    Shift Workers
    ======================================================================
    TO: Yi Qiang
    3rd: Yi Zhang

    Shift Summary
    ======================================================================
    16:00 Shift starts with production
    20:15 Power glitch happened, target HPH went crazy, move to Al dummy
    target to take data.
    20:30 Target HPH recovered, temperature getting normal
    22:20-> ROC5 got lost then several ROCs later on. After some attemps
    (alse paged Bryan Moffit), they were all rebooted. However,
    when we started a run, ROC9 or ROC10 would always get
    disconnected. It is found out to be a CAMAC crater which was
    accidentally rebooted and wrong threshold was set on the
    Bigbite Detector so that ROC9/10 suffered from a buffer
    overflow due the very high rate. A controlled access was
    suggested by Bryan to reboot them.
    23:50 Controlled access to reboot ROC9 and 10.

    Sun Summary
    ======================================================================
    Left HRS + BigBite Production:
    3631 15:47-16:05 LH2 5uA 1M
    3632 16:09-16:27 LH2 5uA 1M
    3633 16:28-16:47 LH2 5uA 1M
    3634 16:47-17:10 LH2 5uA 1M
    3635 17:11-17:34 LH2 5uA 1M
    3636 17:34-17:54 LH2 5uA 1M
    3637 17:54-18:15 LH2 5uA 1M
    3638 18:15-18:33 LH2 5uA 1M
    3639 18:34-18:53 LH2 5uA 1M
    3640 18:54-19:19 LH2 5uA 1M
    3641 19:20-19:37 LH2 5uA 1M
    3642 19:38-19:56 LH2 5uA 1M
    3643 19:57-20:15 LH2 5uA 1M

    3644 20:37-21:21 Al 5uA 1M
    3645 21:22-22:22 Al 5uA 640k

    3646-3648 Junk

    Right HRS
    20732 14:29-17:15 LH2 5uA 277k
    20733 17:16-19:58 LH2 5uA 274k
    20734 19:58-20:15 LH2 5uA 29k
    20735 20:38-22:13 Al 5uA 144k
    20736 Junk
    20737 22:50-00:20 LH2 5uA 129k